What was the return on a 15-year, 7.85 %, $250,000 coupon bond that was purchased one year ago for $245,000

    and just sold for par? Show all work.
Answer #1

Coupon payment = 7.85% * 250000 = 19625

Total return in one year = 250000 + 19625 - 245000 = 24625

Current yield = 24625 / 245000 * 100 = 10.05%